# landmarker: a Toolkit for Anatomical Landmark Localization in 2D/3D Images

### v1

@@ Abstract
Repository: https://github.com/predict-idlab/landmarker.
@@ 1 Motivation and significance
Landmark (or keypoint) localization in 2D/3D images is a fundamental challenge in computer vision, crucial for tasks such as pose estimation (Luvizon et al., 2018; Tompson et al., 2014; Newell et al., 2016), face alignment (Zhou et al., 2023; Kumar et al., 2019), robotic manipulation (Ziegler et al., 2020), surgical procedures (Sharan et al., 2021), and various image-based medical diagnoses (de Queiroz Tavares Borges Mesquita et al., 2023; Ryu et al., 2022; Pei et al., 2023; Payer et al., 2019; Thaler et al., 2021). The effectiveness of different approaches is often domain and task-dependent, necessitating extensive experimentation for new applications. Popular methods, such as heatmap regression, involve numerous preprocessing and post-processing steps. The latter, such as sub-pixel accuracy methods, can significantly improve the precision needed in medical applications but nevertheless are frequently overlooked.
As existing packages such as OpenPose (Cao et al., 2021) and Ultralytics (Jocher et al., 2023) are primarily designed for pose estimation, they lack the modularity needed for medical applications. While MMPose (MMPose Contributors, 2020) offers modular components for various tasks, it is still tailored toward pose estimation and does not easily integrate with other frameworks or handle specific medical imaging requirements.
Our Python package, landmarker [^1] https://github.com/predict-idlab/landmarker, addresses the need for precise anatomical landmark localization in medical images, which is critical for diagnostics and therapeutic procedures in different specialties, such as orthodontics, maxillofacial surgery, and orthopedics. It is a flexible PyTorch-based toolkit designed explicitly for anatomical landmark localization. This toolkit accelerates the development of algorithms and enhances the accuracy of landmark identification in medical images, thereby improving diagnostic and treatment outcomes. Its modular and adaptable framework allows researchers and practitioners to implement state-of-the-art algorithms tailored to their specific datasets.
@@ 2 Background: landmark localization
There are two primary methodologies in landmark localization (see Figure 1): heatmap regression and coordinate regression. Coordinate regression (Fard and Mahoor, 2022) involves learning a direct relationship between the image and the landmark locations. However, the introduction of heatmap regression (Tompson et al., 2014), which generally yields better performance, has shifted the research focus. Heatmap regression predicts a heatmap that assigns a likelihood or probability to each (downsized) image pixel. Early heatmap regression methods, termed direct heatmap regression, assume a parametric distribution, typically a bivariate Gaussian or Laplacian distribution. Direct heatmap regression can be further classified into static and adaptive approaches. In static heatmap regression, the heatmap distribution parameters are fixed hyperparameters during training. In adaptive heatmap regression, these parameters are adjusted during training, either by treating them as learnable model parameters (Payer et al., 2019, 2020; Thaler et al., 2021) or by using scheduling methods that modify the parameters based on specific evaluation metrics (Teixeira et al., 2019). Another variant of static heatmap regression involves training a model on one-hot heatmap images (masks), essentially parameterizing the heatmap as a categorical distribution and transforming the task into a multi-class classification problem where each pixel represents a class (McCouat and Voiculescu, 2022).
Recent advancements on heatmap regression have seen the rise of fully convolutional neural networks (CNNs) with differentiable decoding operations, such as the soft-argmax operation (Luvizon et al., 2018). These methods, prominent in facial landmark localization, use the heatmap generation layer as an intermediate layer and the decoding operation as the final layer, optimizing the network with a loss function that compares the decoded coordinates to the ground-truth coordinates. Some of these methods account for the ambiguity in landmark ground truth by incorporating uncertainty-aware loss functions, such as the Gaussian log-likelihood loss (Kumar et al., 2019), or other custom loss functions (Kumar et al., 2019, 2020; Zhou et al., 2023).
In the medical domain, landmark localization often employs two-stage approaches (Jiang et al., 2022; Song et al., 2020; Zhong et al., 2019). The first stage localizes landmarks on a low-resolution image, and the second stage refines this localization on a high-resolution patch, which is a region of interest based on the initial predictions. In both stages, one of the previously described approaches can be used.
@@ 3 Software description
landmarker is a Python package that leverages PyTorch (Paszke et al., 2019) deep learning framework and MONAI (Cardoso et al., 2022), a PyTorch-based deep learning framework for healthcare imaging for handling medical image files and transformations.
@@ 3.1.1 Data handling and preprocessing
These datasets support various image formats (NIfTI, DICOM, PNG, JPG, BMP, NPY/NPZ) through integration with the MONAI library (Cardoso et al., 2022). The module can handle single-class, multi-class single-instance, and multi-class multi-instance landmark problems. Utility functions are provided to transform common annotation formats (e.g., labelme (Wada, 2024)) into the required format.
The package also includes the datasets module, which contains functions to import benchmark datasets, such as the ISBI2015 (Wang et al., 2016) dataset directly. Preprocessing and data augmentation capabilities are available, leveraging MONAI’s transformations for 2D and 3D data. The package handles affine landmarks transformations through the transforms module.
@@ 3.1.2.2 Heatmap decoding
The argmax operation is the simplest decoder operation (Tompson et al., 2014). The operation takes the maximum value of the heatmap $H_{i}$,
@@ 3.1.3 Models and losses
While landmarker is compatible with any PyTorch model or loss function, it includes implementations of successful approaches from the literature, such as the spatial configuration network (Payer et al., 2019), which is effective for positionally consistent medical images like cephalograms. Also, multiple heatmap regression loss functions are implemented from successful approaches (Thaler et al., 2021; Nibali et al., 2018; Kumar et al., 2019; Feng et al., 2018; Wang et al., 2019; Zhou et al., 2023).
@@ 4 Illustrative examples
With the LandmarkDataset and HeatmapGenerator adequately set up, we can now train a model to output likelihood heatmaps. In Listing 5, we train a SpatialConfigurationNetwork (Payer et al., 2019; Thaler et al., 2021) and parameterization of the Gaussian heatmap (Payer et al., 2020; Thaler et al., 2021).
@@
Figure 4: Example of results of running the prediction_inspection_plot on the ISBI2015 dataset.
@@ 5 Impact and conclusions
In conclusion, landmarker represents a powerful tool for advancing the field of anatomical landmark localization in medical imaging. Its flexibility and ease of use make it a helpful resource for researchers and engineers.
@@
Jef Jonkers is funded by the Research Foundation Flanders (FWO, Ref. 1S11525N).

### v2

@@ Abstract
Code: https://github.com/predict-idlab/landmarker.
Published [^1] Published in SoftwareX: https://doi.org/10.1016/j.softx.2025.102165
@@ 1 Motivation and significance
Landmark (or keypoint) localization in 2D/3D images is a fundamental challenge in computer vision, crucial for tasks such as pose estimation Luvizon et al. (2018); Tompson et al. (2014); Newell et al. (2016), face alignment Zhou et al. (2023); Kumar et al. (2019), robotic manipulation Ziegler et al. (2020), surgical procedures Sharan et al. (2021), image registration (Grewal et al., 2023), various image-based medical diagnoses de Queiroz Tavares Borges Mesquita et al. (2023); Ryu et al. (2022); Pei et al. (2023); Payer et al. (2019); Thaler et al. (2021) and treatment response modeling in longitudinal images (Ou et al., 2015; Cai et al., 2021). The effectiveness of different approaches is often domain and task-dependent, necessitating extensive experimentation for new applications. Popular methods, such as heatmap regression, involve numerous preprocessing and post-processing steps. The latter, such as sub-pixel accuracy methods, can significantly improve the precision needed in medical applications but nevertheless are frequently overlooked.
As existing packages such as OpenPose Cao et al. (2021) and Ultralytics Jocher et al. (2023) are primarily designed for pose estimation, they lack the modularity needed for medical applications. While MMPose MMPose Contributors (2020) offers different components for various tasks, it is still tailored toward pose estimation and does not easily integrate with other frameworks or handle specific medical imaging requirements. Although MMPose (MMPose Contributors, 2020) supports 3D landmark localization, it cannot process 3D input data, a critical requirement for medical imaging applications. Furthermore, while MMPose offers some flexibility in model creation, its pipeline mandates the use of its framework for preprocessing, training, and evaluation, limiting compatibility with deep learning frameworks such as PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019). Additionally, neither pose estimation library supports medical imaging formats (e.g., NIfTI and DICOM) or provides specialized data augmentation techniques to simulate common acquisition artifacts. Table 1 summarizes the key differences between these existing packages and the proposed landmark detection package within the medical imaging context.
3D Medical
Package Custom Keypoints Images Modular Format Aug.
ultralytics (Jocher et al., 2023)
MMPose (MMPose Contributors, 2020)
landmarker
Table 1: Comparison of Python packages for keypoint/landmark localization.
Our Python package, landmarker, addresses the need for precise anatomical landmark localization in medical images, which is critical for diagnostics and therapeutic procedures in different specialties, such as orthodontics, maxillofacial surgery, and orthopedics. It is a flexible PyTorch-based toolkit designed explicitly for training, modeling, and evaluating anatomical landmark localization problems. This toolkit accelerates the development of algorithms and enhances the accuracy of landmark identification in medical images, thereby improving diagnostic and treatment outcomes. Its modular and adaptable framework allows researchers and practitioners to implement state-of-the-art algorithms tailored to their specific datasets.
@@ 2 Background: landmark localization
There are two primary methodologies in landmark localization (see Figure 1): heatmap regression and coordinate regression. Coordinate regression (Fard and Mahoor, 2022) involves learning a direct relationship between the image and the landmark locations. However, the introduction of heatmap regression Tompson et al. (2014), which generally yields better performance, has shifted the research focus. Heatmap regression
@@
predicts a heatmap that assigns a likelihood or probability to each (downsized) image pixel. Early heatmap regression methods, termed direct heatmap regression, assume a parametric distribution, typically a bivariate Gaussian or Laplacian distribution. Direct heatmap regression can be further classified into static and adaptive approaches. In static heatmap regression, the heatmap distribution parameters are fixed hyperparameters during training. In adaptive heatmap regression, these parameters are adjusted during training, either by treating them as learnable model parameters (Payer et al., 2019, 2020; Thaler et al., 2021) or by using scheduling methods that modify the parameters based on specific evaluation metrics (Teixeira et al., 2019). Another variant of static heatmap regression involves training a model on one-hot heatmap images (masks), essentially parameterizing the heatmap as a categorical distribution and transforming the task into a multi-class classification problem where each pixel represents a class (McCouat and Voiculescu, 2022).
Recent advancements on heatmap regression have seen the rise of fully convolutional neural networks (CNNs) with differentiable decoding operations, such as the soft-argmax operation Luvizon et al. (2018). These methods, prominent in facial landmark localization, use the heatmap generation layer as an intermediate layer and the decoding operation as the final layer, optimizing the network with a loss function that compares the decoded coordinates to the ground-truth coordinates. Some of these methods account for the ambiguity in landmark ground truth by incorporating uncertainty-aware loss functions, such as the Gaussian log-likelihood loss (Kumar et al., 2019), or other custom loss functions Kumar et al. (2019, 2020); Zhou et al. (2023).
In the medical field, landmark localization often utilizes two-stage approaches (Jiang et al., 2022; Song et al., 2020; Zhong et al., 2019; Nguyen et al., 2020). The first stage identifies landmarks on a low-resolution image, while the second stage refines this localization using a high-resolution patch, which is a region of interest based on the initial predictions. One of the previously described approaches can be applied in both stages. Another technique, also used in medical imaging, involves segmenting or contouring specific anatomical structures and inferring the landmarks from the resulting masks or contours (Gooßen et al., 2011; Pei et al., 2021).
@@ 3 Software description
landmarker is a Python package that leverages PyTorch Paszke et al. (2019) deep learning framework and MONAI Cardoso et al. (2022), a PyTorch-based deep learning framework for healthcare imaging for handling medical image files and transformations.
@@ 3.1.1 Data handling and preprocessing
These datasets support various image formats (NIfTI, DICOM, PNG, JPG, BMP, NPY/NPZ) through integration with the MONAI library (Cardoso et al., 2022). The module can handle single-class, multi-class single-instance, and multi-class multi-instance landmark problems. Utility functions are provided to transform common annotation formats (e.g., labelme Wada (2024)) into the required format.
The package also includes the datasets module, which contains functions to import benchmark datasets, such as the ISBI2015 Wang et al. (2016) dataset directly. Preprocessing and data augmentation capabilities are available, leveraging MONAI’s transformations for 2D and 3D data. The package handles affine landmarks transformations through the transforms module.
@@ 3.1.2.2 Heatmap decoding
The argmax operation is the simplest decoder operation Tompson et al. (2014). The operation takes the maximum value of the heatmap $H_{i}$,
@@ 3.1.3 Models and losses
While landmarker is compatible with any PyTorch model or loss function, it includes implementations of successful approaches from the literature, such as the spatial configuration network Payer et al. (2019), which is effective for positionally consistent medical images like cephalograms. Also, multiple heatmap regression loss functions are implemented from successful approaches (Thaler et al., 2021; Nibali et al., 2018; Kumar et al., 2019; Feng et al., 2018; Wang et al., 2019; Zhou et al., 2023).
@@ 4 Illustrative examples
With the LandmarkDataset and HeatmapGenerator adequately set up, we can now train a model to output likelihood heatmaps. In Listing 5, we train a SpatialConfigurationNetwork Payer et al. (2019); Thaler et al. (2021) and parameterization of the Gaussian heatmap Payer et al. (2020); Thaler et al. (2021).
@@
(a) ISBI2015 dataset
(b) Pelvis X-rays
(c) Endoscopic images
Figure 4: Examples of landmark predictions using the prediction_inspection_plot tool for different datasets: (a) Endoscopic images, (b) ISBI2015, and (c) Pelvis X-rays.
4.1 Benchmarks
To evaluate our package capabilities and its model’s landmark localization precision quantitatively, we conducted two benchmark experiments on relatively recent datasets that showcase its performance on both 2D and 3D medical imaging tasks. For the 2D benchmark, we used pelvis x-ray images from the Osteoarthritis Initiative (Pei et al., 2023). In this experiment, our spatial configuration network (SCN) with adaptive heatmaps—implemented in landmarker, was compared against a state-of-the-art U-Net with Attention approach (Pei et al., 2023). As shown in Table 2, our method achieved a lower average position error (PE) and higher success detection rates (SDR) across multiple thresholds, clearly demonstrating the capabilities of the package as a benchmarking tool.
Table 2: Landmark localization results of spatial configuration network (SCN) with adaptive heatmaps (Payer et al., 2020; Thaler et al., 2021) (implemented in landmarker) on the pelvis x-rays (2D) of the osteoarthritis initiative dataset (Pei et al., 2023) compared against the literature. The best results are highlighted in bold. A 5-fold validation approach, as suggested by Pei et al. (Pei et al., 2023), obtains the results.
Model name PE (mm) SDR (%)
1 mm 2 mm 3 mm 4 mm
U-Net w/ Attention (Pei et al., 2023)
3.14 8.65% 33.89% 59.25% 78.61%
SCN (Payer et al., 2020; Thaler et al., 2021)
1.61 33.25% 74.38% 90.96% 95.85%
In addition to the 2D evaluation, we benchmarked our model on 3D images using CT scans of the skull (He et al., 2024), where the task was to annotate mandibular landmarks automatically. The results, presented in Table 3, include both validation and test metrics, with our one-hot ensemble (Jonkers et al., 2025) method outperforming the competitive baseline regarding PE and SDR. These results confirm our model’s precision and highlight its flexibility in handling diverse anatomical landmarking applications across different imaging modalities.
Table 3: Landmark localization results of one hot ensemble (Jonkers et al., 2025) (implemented in landmarker) on the MML (3D) dataset compared against the literature. This benchmark dataset is the same as the data subset, only with complete landmarks, used as a benchmark in He et al. (He et al., 2024). The best results are highlighted in bold.
Model name Validation Test
PE (mm) SDR (%) PE (mm) SDR (%)
2 mm 2.5 mm 3 mm 4 mm 2 mm 2.5 mm 3 mm 4 mm
Pruning-ResUNet3D (He et al., 2024)
1.82 73.21% 82.14% 88.93% 94.76% 1.96 70.03% 79.97% 86.10% 92.73%
One hot ensemble (Jonkers et al., 2025)
1.60 77.81% 87.76% 89.92% 93.37% 1.39 81.67% 91.31% 93.33% 96.31%
@@ 5 Impact and conclusions
In conclusion, landmarker represents a powerful tool for advancing the field of anatomical landmark localization in medical imaging. Its flexibility and ease of use make it a helpful resource for researchers and engineers. We plan to incorporate future developments and improvements, particularly in uncertainty quantification, which is critical for enhancing the reliability and robustness of landmark localization in medical applications.
CRedit Authorship Contribution Statement
Jef Jonkers: Conceptualization, Methodology, Software, Validation, Visualization, Writing - Original Draft. Luc Duchateau: Conceptualization, Writing - Review & Editing, Supervision, Funding acquisition. Glenn Van Wallendael: Conceptualization, Writing - Review & Editing, Supervision. Sofie Van Hoecke Conceptualization, Writing - Review & Editing, Supervision, Funding acquisition.
@@
Jef Jonkers is funded by the Research Foundation Flanders (FWO, Ref. 1S11525N). This research was (partially) funded by the Flemish Government (AI Research Program).